Find the directional derivative of (x,y,z) = xsin yz at (1,3,0) in the direction of i+2j-k


1
Expert's answer
2021-08-23T14:49:24-0400
"f(x,y,z)=x\sin yz"

"\nabla f=(\sin yz)i+(xz\cos yz)j+(xy\cos yz)k"


"\nabla f(1,3,0)=0i+0j+3k"

"|i+2j-k|=\sqrt{1^2+2^2+(-1)^2}=\sqrt{6}"

"u=\dfrac{1}{\sqrt{6}}i+\dfrac{2}{\sqrt{6}}j-\dfrac{1}{\sqrt{6}}k"

"D_uf(1,3,0)=(0i+0j+3k)(\dfrac{1}{\sqrt{6}}i+\dfrac{2}{\sqrt{6}}j-\dfrac{1}{\sqrt{6}}k)"

"=-\dfrac{\sqrt{6}}{2}"